OpenStudy (anonymous):

Solve each exponential equation. Give the exact value for x. (a) 6x = 1/36 x = (b) 3^x = 6 x = (c) 8^x = 6.5 x = (d) 4^x = 5

OpenStudy (mayankdevnani):

@monicaed1971 do you try your question ?

OpenStudy (anonymous):

I am not sure how to do it.

OpenStudy (mayankdevnani):

okk...

OpenStudy (mayankdevnani):

1)\[\large \bf 6x=\frac{1}{36}\] divide both sides by 6,we get, \[\large \bf x=\frac{1}{36 \times 6}=\frac{1}{216}\]

OpenStudy (mayankdevnani):

understood ?

OpenStudy (anonymous):

yes, thank you very much

OpenStudy (anonymous):

How is be done it is a little different and has and exponent of x?

OpenStudy (mayankdevnani):

2) \[\large \bf 3^x=6\] we can use LOG to solve this problem.

OpenStudy (mayankdevnani):

so do you know what is LOG ?

OpenStudy (anonymous):

yes, I am familiar with log.

OpenStudy (mayankdevnani):

okk

OpenStudy (mayankdevnani):

2) \[\large \bf 3^x=6\] \[\large \bf \log{3^x}=\log 6\] we use log to the base 10, then, \[\large \bf xlog3=\log6\] \[\large \bf x=\frac{\log6}{\log3}\] \[\large \bf where,\log3=0.47712125472 ~and~\log6=0.77815125038\] plug the values and get your answer

OpenStudy (anonymous):

1.630929753562667